# On degenerate metric resolvent: well-definedness and new convergence analysis

### v3

@@ 3.1 Main result
Combining Lemma 2.10-(4) and Theorem 3.4, we claim that $\mathcal{T}$ has full domain, if and only if $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone. What remains to show is the equivalence of maximal monotonicity between $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ and $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$.
First, if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have:
$(u,x)\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v,y)\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u-v|x-y\rangle\geq 0.$
Since $u,x,v,y\in\mathsf{ran}\mathcal{Q}$, let $u=\sqrt{\mathcal{Q}}u^{\prime}$ and $v=\sqrt{\mathcal{Q}}v^{\prime}$ for $u^{\prime},v^{\prime}\in\mathsf{ran}\mathcal{Q}$. Then, $x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}(u)=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}u=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}u^{\prime}$, and similarly, $y\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v^{\prime}$. Then, we obtain
$(u^{\prime},x)\in\mathsf{gra}\Big{(}\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\Longleftrightarrow\forall(v^{\prime},y)\in\mathsf{gra}\Big{(}\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)},\ \langle\sqrt{\mathcal{Q}}(u^{\prime}-v^{\prime})|x-y\rangle\geq 0.$
Let $x^{\prime}=\sqrt{\mathcal{Q}}x$ and $y^{\prime}=\sqrt{\mathcal{Q}}y$, then it becomes
$(u^{\prime},x^{\prime})\in\mathsf{gra}\Big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\Longleftrightarrow\forall(v^{\prime},y^{\prime})\in\mathsf{gra}\Big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)},\ \langle u^{\prime}-v^{\prime}|x^{\prime}-y^{\prime}\rangle\geq 0.$
This shows that $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
Conversely, if $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have:
$(u,x)\in\mathsf{gra}(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v,y)\in\mathsf{gra}(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u-v|\sqrt{\mathcal{Q}}x-\sqrt{\mathcal{Q}}y\rangle\geq 0.$
Let $x^{\prime}=\sqrt{\mathcal{Q}}x$, $y^{\prime}=\sqrt{\mathcal{Q}}y$, $u^{\prime}=\sqrt{\mathcal{Q}}u$, $v^{\prime}=\sqrt{\mathcal{Q}}v$, we have:
$(u^{\prime},x^{\prime})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v^{\prime},y^{\prime})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u^{\prime}-v^{\prime}|x^{\prime}-y^{\prime}\rangle\geq 0.$
This shows the maximal monotonicity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$.
@@
According to Lemma 3.2, Assumption 2 has already been implied by the maximal monotonicity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$. Thus, there is no need to put Assumption 2 in Theorem 3.1.
@@ 3.2 Sufficient condition
The next question regarding Theorem 3.1 naturally arises: what is the condition on $\mathcal{A}$ and $\mathcal{Q}$, for which $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone? The following important corollary gives a sufficient condition.
@@
Under Assumption 1, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if the following conditions are fulfilled:
@@
(ii) $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$ [^1] The definition of strong relative interior $(\mathsf{sri})$ is given by [4, Definition 6.9]..
In particular, if $\mathcal{Q}$ is non-degenerate, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if and only if $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone.
@@
In view of [4, Corollary 25.6 or Proposition 25.41-(iv)].
@@
$\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ (or $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$): $\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
@@
(ii) $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$.
@@
In addition, without Assumption 4-(ii), we claim that $\mathcal{T}$almost’ has full domain.
@@
Under Assumptions 1 and 4-(i), $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})\supseteq\mathsf{ran}\mathcal{Q}$.
@@
First, $\mathcal{Q}$ is maximally monotone by [4, Corollary 20.28], and $3^{*}$ monotone by Fact 1-(iv) and [4, Example 25.17]. Then, $\mathcal{A}+\mathcal{Q}$ is maximally monotone by [4, Corollary 25.5].
Then, noting $\mathsf{dom}\mathcal{A}\subseteq\mathsf{dom}\mathcal{Q}=\mathcal{H}$, [4, Theorem 25.24] yields
$\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})=\overline{\mathsf{ran}\mathcal{A}+\mathsf{ran}\mathcal{Q}}\supseteq\overline{0+\mathsf{ran}\mathcal{Q}}=\overline{\mathsf{ran}}\mathcal{Q}=\mathsf{ran}\mathcal{Q},$
where the last equality is due to the closedness of $\mathsf{ran}\mathcal{Q}$.
The above result always holds, without Assumption 4-(ii). One can check Example 3.12, where $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\nsupseteq\mathsf{ran}\mathcal{Q}$ but $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})=\mathbb{R}\times\mathbb{R}_{+}\supset\mathbb{R}\times\{0\}=\mathsf{ran}\mathcal{Q}$. Unfortunately, it is generally difficult to investigate the relation between $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})$ and $\mathsf{ran}(\mathcal{A}+\mathcal{Q})$. We believe there are some connections between Assumption 4-(ii) and $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})$. This is left to future work.
@@ 3.3 Examples
We list several examples to show
@@
(2) $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ may not be maximally monotone, even if $\mathcal{A}$ is maximally monotone;
@@
It is easy to see that $\mathsf{ran}\mathcal{A}=[-1,1]\times\{0\}$, $\mathsf{ran}\mathcal{Q}=\{0\}\times\mathbb{R}$, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=[-1,1]\times\mathbb{R}\neq\mathcal{H}=\mathbb{R}^{2}$, but it already suffices to guarantee that $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\supset\mathsf{ran}\mathcal{Q}$. One further has $\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q}=[-1,1]\times\mathbb{R}$, which contains 0 as a relative interior.
Is $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ maximally monotone? First, noting that $\mathsf{dom}\mathcal{A}^{-1}=\mathsf{ran}\mathcal{A}=[-1,1]\times\{0\}$, $\mathcal{A}^{-1}$ is given as:
$\mathcal{A}^{-1}:(x,y)\mapsto\left\{\begin{array}[]{ll}\mathbb{R}_{+}\times\mathbb{R},&\textrm{if\ }x=1,\ y=0;\\ \mathbb{R}_{-}\times\mathbb{R},&\textrm{if\ }x=-1,\ y=0;\\ \{0\}\times\mathbb{R},&\textrm{if\ }x\in(-1,1),\ y=0.\\ \end{array}\right.$
where +=$\mathbb{R}_{+}=\{x\in\mathbb{R}:x\geq 0\}$, =$\mathbb{R}_{-}=\{x\in\mathbb{R}:x\leq 0\}$.
We now check whether $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})=\{0\}\times\mathcal{R}=\mathsf{ran}\mathcal{Q}$. Indeed, we develop:
$\mathsf{ran}\Big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}$ $\supseteq$ $\Big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\begin{bmatrix}0\\ 0\end{bmatrix}$
$=$ $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}0\\ 0\end{bmatrix}=\sqrt{\mathcal{Q}}(\{0\}\times\mathbb{R})$
$=$ $\{0\}\times\mathbb{R}=\mathsf{ran}\mathcal{Q},$
which shows $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone by restricted Minty’s theorem.
@@
In [10, Remark 2.2], the authors proposed an interesting counter-example that deserves particular treatment. The problem is
@@
We then deduce that $\mathsf{ran}\mathcal{A}=[-1,0]\times\mathbb{R}_{+}$, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=(-\infty,-1]\times(0,+\infty)\cup(-1,+\infty]\times[0,+\infty)=\mathbb{R}\times[0,+\infty)\backslash(-\infty,-1]\times\{0\}$. Clearly, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\nsupseteq\mathsf{ran}\mathcal{Q}=\mathbb{R}\times\{0\}$. Note that $(\mathcal{A}+\mathcal{Q})^{-1}(x,0)=\varnothing$ for $x\leq-1$, though $(x,0)\in\mathsf{ran}\mathcal{Q}$.
Since $\mathcal{T}$ is not of full domian, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ must not be maximally monotone by Theorem 3.1. Let us check it. First, we have $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}:(x,0)\mapsto\mathbb{R}_{++}\times\{0\}$, $\forall x\in[-1,0]$, where ++=$\mathbb{R}_{++}=\{x\in\mathbb{R}:x>0\}$. Then, $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})=(-1,+\infty)\times\{0\}\neq\mathbb{R}\times\{0\}$, and $(-\infty,-1]\nsubseteq\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})$. Thus, by Theorem 3.4, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is not maximally monotone.
Finally, it is easy to check that Assumption 4-(ii) is not satisfied. Indeed, $\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A}=\mathbb{R}\times(-\infty,0]$, and thus, $0\notin\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$. This shows that Assumption 4 is tight.

### v4

@@ 3.1 Main result
Combining Lemma 2.11-(4) and Theorem 3.4, we claim that $\mathcal{T}$ has full domain, if and only if $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone. What remains to show is the equivalence of maximal monotonicity between $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ and $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathrm{ran}\mathcal{Q}}$.
First, if $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathrm{ran}\mathcal{Q}\times\mathrm{ran}\mathcal{Q}$, we have:
$(u,x)\in\mathrm{gra}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}})\Longleftrightarrow\forall(v,y)\in\mathrm{gra}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}),\ \langle u-v|x-y\rangle\geq 0.$ (8)
Let $u=\sqrt{\mathcal{Q}}u^{\prime}$ and $v=\sqrt{\mathcal{Q}}v^{\prime}$ for $u^{\prime},v^{\prime}\in\mathrm{ran}\mathcal{Q}$. Then, $x\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}u^{\prime}$ and $y\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v^{\prime}$. Further, let $x^{\prime}=\sqrt{\mathcal{Q}}x$ and $y^{\prime}=\sqrt{\mathcal{Q}}y$, then (8) becomes
$(u^{\prime},x^{\prime})\in\mathrm{gra}\Big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}\Big{)}\Longleftrightarrow\forall(v^{\prime},y^{\prime})\in\mathrm{gra}\Big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}\Big{)},\ \langle u^{\prime}-v^{\prime}|x^{\prime}-y^{\prime}\rangle\geq 0.$
This shows that $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone.
The converse statement can be shown in a similar manner.
@@
The monotonicity of $\mathrm{gra}\mathcal{A}\cap(\mathcal{H}\times\mathrm{ran}\mathcal{Q})$ is compulsory in Theorems 3.1 and 3.4. There are many possibilities for $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$ to satisfy $\mathrm{ran}(\mathcal{I}+\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}})=\mathrm{ran}\mathcal{Q}$ beyond the restricted monotonicity of $\mathcal{A}$ within ×$\mathcal{H}\times\mathrm{ran}\mathcal{Q}$. This sufficient and necessary condition holds only under the prerequisite of the resticted monotonicity. This is very similar to the well-known Minty’s theorem [5, Theorem 21.1], where the conclusion holds only when $\mathcal{A}$ is assumed to be monotone.
@@ 3.2 Sufficient condition
The next question regarding Theorem 3.1 naturally arises: what is the condition on $\mathcal{A}$ and $\mathcal{Q}$, for which $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathrm{ran}\mathcal{Q}}:\mathrm{ran}\mathcal{Q}\mapsto 2^{\mathrm{ran}\mathcal{Q}}$ is maximally monotone? The following important corollary gives a sufficient condition.
@@
Under Assumption 1, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone, if the following conditions are fulfilled:
@@
(ii) $0\in\mathrm{sri}(\mathrm{ran}\mathcal{Q}-\mathrm{ran}\mathcal{A})$, where the strong relative interior $(\mathrm{sri})$ is defined in [5, Definition 6.9].
In particular, if $\mathcal{Q}$ is non-degenerate, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone, if and only if $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone.
@@
In view of [5, Corollary 25.6 or Proposition 25.41-(iv)].
@@
$\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ (or $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathrm{ran}\mathcal{Q}}$): $\mathrm{ran}\mathcal{Q}\mapsto 2^{\mathrm{ran}\mathcal{Q}}$ is maximally monotone.
@@
(ii) $0\in\mathrm{sri}(\mathrm{ran}\mathcal{Q}-\mathrm{ran}\mathcal{A})$.
By Corollary 3.7, both assumptions are equivalent for non-degenerate case, since Assumption 4-(ii) is automatically satisfied. Indeed, $\mathrm{ran}\mathcal{Q}-\mathrm{ran}\mathcal{A}\supseteq\mathrm{ran}\mathcal{Q}-0=\mathrm{ran}\mathcal{Q}=\mathcal{H}$, which obviously contains 0 in its interior (and also strong relative interior).
@@
In addition, without Assumption 4-(ii), the next proposition claims that $\mathcal{T}$almost’ has full domain.
@@
Under Assumptions 1, 2 and 4-(i), $\overline{\mathrm{ran}}(\mathcal{A}+\mathcal{Q})\supseteq\mathrm{ran}\mathcal{Q}$.
@@
First, $\mathcal{Q}$ is maximally monotone by [5, Corollary 20.28], and $3^{*}$ monotone by Fact 1-(iv) and [5, Example 25.17]. Then, $\mathcal{A}+\mathcal{Q}$ is maximally monotone by [5, Corollary 25.5]. Then, noting $\mathrm{dom}\mathcal{A}\subseteq\mathrm{dom}\mathcal{Q}=\mathcal{H}$, [5, Theorem 25.24] yields
$\overline{\mathrm{ran}}(\mathcal{A}+\mathcal{Q})=\overline{\mathrm{ran}\mathcal{A}+\mathrm{ran}\mathcal{Q}}\supseteq\overline{0+\mathrm{ran}\mathcal{Q}}=\overline{\mathrm{ran}}\mathcal{Q}=\mathrm{ran}\mathcal{Q},$
which is due to $0\in\mathrm{ran}\mathcal{A}$ (see Assumption 2-(i)) and the closedness of $\mathrm{ran}\mathcal{Q}$.
The above result always holds, without Assumption 4-(ii). One can check Example 3.12 (shown later), where $\mathrm{ran}(\mathcal{A}+\mathcal{Q})\nsupseteq\mathrm{ran}\mathcal{Q}$ but $\overline{\mathrm{ran}}(\mathcal{A}+\mathcal{Q})=\mathbb{R}\times\mathbb{R}_{+}\supset\mathbb{R}\times\{0\}=\mathrm{ran}\mathcal{Q}$. Unfortunately, it is generally difficult to investigate the relation between $\overline{\mathrm{ran}}(\mathcal{A}+\mathcal{Q})$ and $\mathrm{ran}(\mathcal{A}+\mathcal{Q})$. We believe there are some underlying connections between Assumption 4-(ii) and $\overline{\mathrm{ran}}(\mathcal{A}+\mathcal{Q})$. This is left to future work.
@@ 3.3 Examples
We list several examples to show (in degenerate case)
@@
(2) $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ may not be maximally monotone, even if $\mathcal{A}$ is maximally monotone;
@@
It is easy to see that $\mathrm{ran}\mathcal{A}=[-1,1]\times\{0\}$, $\mathrm{ran}\mathcal{Q}=\{0\}\times\mathbb{R}$, $\mathrm{ran}(\mathcal{A}+\mathcal{Q})=[-1,1]\times\mathbb{R}\neq\mathcal{H}=\mathbb{R}^{2}$, but it already suffices to guarantee that $\mathrm{ran}(\mathcal{A}+\mathcal{Q})\supset\mathrm{ran}\mathcal{Q}$. One further has $\mathrm{ran}\mathcal{A}-\mathrm{ran}\mathcal{Q}=[-1,1]\times\mathbb{R}$, which contains 0 as a relative interior. Thus, Corollary 3.7 concludes that $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone.
Let us verify it in another way. First, noting that $\mathrm{dom}\mathcal{A}^{-1}=\mathrm{ran}\mathcal{A}=[-1,1]\times\{0\}$, $\mathcal{A}^{-1}$ is given as:
$\mathcal{A}^{-1}:(x,y)\mapsto\left\{\begin{array}[]{ll}[0,+\infty)\times\mathbb{R},&\textrm{if\ }(x,y)=(1,0);\\ (-\infty,0]\times\mathbb{R},&\textrm{if\ }(x,y)=(-1,0);\\ \{0\}\times\mathbb{R},&\textrm{if\ }(x,y)\in(-1,1)\times\{0\}.\\ \end{array}\right.$
We now check whether $\mathrm{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}})=\{0\}\times\mathbb{R}=\mathrm{ran}\mathcal{Q}$. Indeed, we develop: $\mathrm{ran}\Big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}\Big{)}\supseteq\Big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}\Big{)}\begin{bmatrix}0\\ 0\end{bmatrix}=\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}0\\ 0\end{bmatrix}=\{0\}\times\mathbb{R}=\mathrm{ran}\mathcal{Q}$, which shows $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ is maximally monotone by restricted Minty’s theorem.
@@
In [13, Remark 2.2], the authors proposed an interesting counter-example that deserves particular treatment. The problem is
@@
We then deduce that $\mathrm{ran}\mathcal{A}=[-1,0]\times\mathbb{R}_{+}$, $\mathrm{ran}(\mathcal{A}+\mathcal{Q})=(-\infty,-1]\times(0,+\infty)\cup(-1,+\infty]\times[0,+\infty)=\mathbb{R}\times[0,+\infty)\backslash(-\infty,-1]\times\{0\}$. Clearly, $\mathrm{ran}(\mathcal{A}+\mathcal{Q})\nsupseteq\mathrm{ran}\mathcal{Q}=\mathbb{R}\times\{0\}$. Note that $(\mathcal{A}+\mathcal{Q})^{-1}(x,0)=\varnothing$ for $x\leq-1$, though $(x,0)\in\mathrm{ran}\mathcal{Q}$.
Since $\mathcal{T}$ is not of full domian, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ must not be maximally monotone by Theorem 3.1. Let us check it. First, we have $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}:(x,0)\mapsto(0,+\infty)\times\{0\}$, $\forall x\in[-1,0]$. Then, $\mathrm{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}})=(-1,+\infty)\times\{0\}\neq\mathbb{R}\times\{0\}$, and $(-\infty,-1]\nsubseteq\mathrm{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}})$. Thus, by Theorem 3.4, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ is not maximally monotone.
Finally, it is easy to check that Assumption 4-(ii) is not satisfied. Indeed, $\mathrm{ran}\mathcal{Q}-\mathrm{ran}\mathcal{A}=\mathbb{R}\times(-\infty,0]$, and thus, $0\notin\mathrm{sri}(\mathrm{ran}\mathcal{A}-\mathrm{ran}\mathcal{Q})$. This shows that Assumption 4 is tight.